OpenStudy (anonymous):

HELP PRECALC GIVING MEDALS AND BECOMING A FAN 1.)Find an explicit rule for the nth term of the sequence. -4, -8, -16, -32 an = -4 • 2^(n - 1) an = 2 • -4^(n + 1) an = 2 • -4^n an = -4 • 2^n 2.) Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively. an = 3 • (-4)^(n + 1) an = 3 • 4^(n - 1) an = 3 • (-4)^(n - 1) an = 3 • 4^n

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you can always check which ones work, but each term in negative, so the number out fron must be negative

OpenStudy (anonymous):

For Question one this is answer: an = -4 • 2n - 1

OpenStudy (misty1212):

if \(n=1\) you should get \(-4\) and i think only \[-4\times 2^{n-1}\] works

OpenStudy (anonymous):

@misty1212 so is an = -4 • 2^(n - 1) correct?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

thank you both @misty1212 for the second question I know it isn't an = 3 • 4^(n - 1) but i hv no idea what it is

OpenStudy (anonymous):

@freckles

OpenStudy (misty1212):

you are supposed to get something that looks like \[a_n=a_0\times r^{n-1}\]

OpenStudy (misty1212):

\[a_2=-12\\ a_5=768\]

OpenStudy (anonymous):

is it A? @misty1212

OpenStudy (misty1212):

divide and get \[\frac{a_5}{a_2}=\frac{768}{-12}=-64\]

OpenStudy (misty1212):

idk i can't eyeball it, we have to do it

OpenStudy (misty1212):

that means \[r^3=-64\] so \[r=-4\]

OpenStudy (misty1212):

then go with C since it is the one that looks like \[a_0r^{n-1}\] in this case \(r=-4\) and \(a_0=3\) since \[3\times (-4)^{2-1}=3\times -4=-12\]

OpenStudy (anonymous):

thanks @misty1212